OpenStudy (anonymous):

Is the system consistent or inconsistent? 2x + 3y = 5 and 4x + 6y = 19 consistent inconsistent

OpenStudy (anonymous):

If you look at the two equations you can see that for the left hand side, the second is simply double the first! As in, 4x+6y is double 2x+3y. So, if the system were consistent it would follow that the right hand side would be double, but 19 isn't double 5, so the system must be inconsistent!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The system is inconsistent since there is no solution for the system. ----------- Ideas: 2x+3y = 5 is equivalent to 4x+6y = 10. So, 4x+6y = 19 at the same time is impossible.

OpenStudy (anonymous):

rule for inconsistency a1/a2=b1/b2 but not =c1/c2. here a1=2 a2=4 b1=3 b2=6 c1=5 c2=19 so the sytem in inconsistent n indeterminant

OpenStudy (precal):



OpenStudy (precal):

no solution